How would you solve this. I got 0 & -1 but they dont plug in the solution right.
5x^2-5x=0

Answer by stanbon(75887)

5x^2-5x=0
5x(x-1) = 0
x = 0 or x = 1
------------------
Checking x=0
5(0^2)-5(0) = 0
0-0 = 0
--------------
Checking x=1
5(1)^2-5*1 = 0
5-5=0
0=0
========
Cheers,
Stan H.